# Effects of dynamical friction on perturbations for evolving dark energy

### v1

@@ IV Evolution of linear perturbations
Even though small, $\eta$ affects the initial conditions of Eq. (5), similarly to what happens in early dark energy and massive neutrinos scenarios. This happens because the friction term $\eta/H$ is small but not negligible. To elucidate this concept, let us denote with $\eta_{\rm ini}=\eta(a_{\rm ini})/H(a_{\rm ini})\ll 1$ the value to be used at initial time to evaluate the initial conditions. In general, at early time, a given cosmological model is well approximated by an EdS model, so that we can assume $\Omega_{\rm m}(t)\approx 1$ and $H^{\prime}/H\approx-3/2$ and that the growth factor follows a power-law behaviour, $\delta=Aa^{n}$, where $A$ is an arbitrary constant amplitude which can be safely set to one. Plugging the Ansatz for $\delta$ into Eq. (5), we need to solve a second order algebraic equation for the exponent $n$:
@@
whose growing mode solution is, for $\eta_{\rm ini}\ll 1$, $n_{+}=1-2\eta_{\rm ini}/5$. This result is expected from physically considerations: friction slows down the collapse and $n_{+}<1$ when $\eta\neq 0$. We can compare this result with a generic early dark energy model. If we denote with $\Omega_{\rm e}\ll 1$ the amount of early dark energy, we have $\Omega_{\rm m}\approx 1-\Omega_{\rm e}$ and we find for the growing mode $n_{+}=1-3\Omega_{\rm e}/5$. This same expression applies to massive neutrinos, if we replace $\Omega_{\rm e}$ with the fraction of massive neutrinos $f_{\nu}=\Omega_{\nu}/\Omega_{\rm m}$, with $\Omega_{\nu}$ the neutrinos density parameter. Note the similarity of the different physical systems. Also in this case, therefore, structure formation is slowed down. This means that to match the pace of structures in the $\Lambda$CDM model, the initial amplitude of perturbations must be higher. We will discuss this in more detail when investigating the evolution of nonlinear perturbations.
After this discussion, we can now study the evolution of the growth factor. We consider two clustering species, baryons (denoted with the subscript $\mathrm{b}$) and dark matter (denoted with the subscript $\mathrm{c}$), of which only the latter is affected by dynamical friction. The two equations to be solved are
@@
$\delta_{\rm m}=\frac{\Omega_{\rm b}\delta_{\rm b}+\Omega_{\rm c}\delta_{\rm c}}{\Omega_{\rm b}+\Omega_{\rm c}}\,.$ (33)
IV.1 Initial settings
If we consider a single clustering species, we can assume that at early times $\delta=A\,a^{n}$, where both $A$ and $n$ are constant. This leads to a second-order algebraic equation for $n$ whose solution, for $\eta=\eta_{0}H$ and $\eta_{0}\ll 1$, is $n=1-\tfrac{2}{5}\eta_{0}$. We can proceed in a similar way if we consider both dark matter and baryons. Also in this case we can assume a power-law, such that $\delta_{\rm b}=A\,a^{n}$ and $\delta_{\rm c}=B\,a^{n}$, with $A$, $B$ and $n$ constants to determine. We assume that the two perturbations have the same slope but different amplitudes. If we plug this Ansatz in the previous equations, we obtain, taking into account that at early times the background is extremely well represented by an EdS model,
@@ IV.1 Initial settings
$\frac{\delta_{\rm b}}{\delta_{\rm c}}=1+\frac{\eta_{\rm ini}}{n+1/2}\,,$ (35)
which shows that the baryon perturbation amplitude is increased with respect to that of cold dark matter, as baryons are not affected by friction. Also note that for vanishing friction, $\delta_{\rm b}=\delta_{\rm c}=\delta_{\rm m}$, as expected.
Figure 1: Evolution of the growth factor for the $\Lambda$CDM model for the two recipes used in this work to describe cosmic friction. Top panel: dynamical friction as parametrised in Eq. (20). Bottom panel: dissipative pressure friction as parametrised in Eq. (21). Different colours and line styles refer to different values of the free parameters.
Plugging Eq. (35) into the equation for $\delta_{\rm c}$, leads to an equation for the exponent $n$:
@@
Note that, at early times $\Omega_{\rm b}=\Omega_{\rm b,0}/\Omega_{\rm m,0}$, with $\Omega_{\rm m,0}=\Omega_{\rm b,0}+\Omega_{\rm c,0}$.
Figure 2: Evolution of the growth factor for the six dark energy models investigated in this work with dynamical friction. Line styles and colours are as in Fig. 1.
Even though it is possible to rewrite it as a third order equation, it is simpler to solve it as it is, to avoid going through complex radicals. We can, however, use the cubic equation to investigate the behaviour of the initial slope when $\eta_{\rm ini}\ll 1$. The equation to be solved is
@@
If $\eta_{\rm ini}=0$, it is easy to see that the solutions are $n=-\tfrac{1}{2}$, $n=-\tfrac{3}{2}$ and $n=1$, of which only the latter two are the physical ones, corresponding to the decaying and growing mode, respectively. At first order in $\eta_{\rm ini}$ we obtain $n=1-\tfrac{2}{5}\eta_{\rm ini}(1-\Omega_{\rm b})$, where it is understood that $\Omega_{\rm b}$ is evaluated at early times. The term in parenthesis is positive for realistic values of the baryon density, hence, also in this case we can see that friction dampens structure formation. If we set, for simplicity $1-\Omega_{\rm b}\approx 1$, we recover the previous result, showing the excellent agreement between the two approaches.
@@
Since the two equations of motion are different, we can not combine them into a single equation for the variable $\delta_{\rm m}$. Therefore, to highlight the importance of friction and the differences induced by having two species rather than one, we consider the evolution of the linear growth factor $D_{+}$ as a function of redshift $z$ considering either one or two species. As differences among the two different cases are of the order of $0.1\%$, for simplicity, we will consider just the case where both dark matter and baryons are affected by dynamical friction. This will also simplify the discussion when considering clustering dark energy models.
Figure 3: As in Fig. 2, but with dissipative pressure friction.
IV.2 Growth factor
@@
In this Section we present the effects of friction on the evolution of the linear growth factor, using the two recipes described above. We first consider its effect on the $\Lambda$CDM reference model. We then compare it to six different dark energy models, two of them characterised by a constant equation of state, $w_{\rm de}=-0.9$ and $w_{\rm de}=-1.1$ (DE1 and DE2), and four with a linearly time-dependent equation of state, $w_{\rm de}=w_{0}+w_{a}(1-a)$, with $w_{0}$ and $w_{a}$ constant (Chevallier and Polarski, 2001; Linder, 2003). We choose the following combinations of parameters $(w_{0},w_{a})$: $(-0.9,\pm 0.3)$ (DE3 and DE4) and $(-1.1,\pm 0.3)$ (DE5 and DE6).
@@
In Fig. 1 we present the evolution of the linear growth factor, divided by the scale factor, for the $\Lambda$CDM model, showing results for both dynamical friction (upper panel) and dissipative pressure friction (bottom panel). The considerations which we are going to make for this model will apply also to the other models studied in this work.
@@
As $\eta$ opposes the formation of structures, perturbations grow more slowly with respect to a model where the friction term is absent. Hence, because we want the same rate today (all the curves have been normalised to unity today), structures need to growth faster in the past. This explains the differences observed. Note that this behaviour is analogous to $\Lambda$CDM versus EdS. Therefore, the higher is the value of $\eta$, the higher the difference from a model without friction.
@@
For dynamical friction, in general, differences are very small and only appreciable for the highest value chosen, $\eta_{0}=10^{-2}$. Effects of dissipative pressure friction follow the same trend. Dissipative pressure friction behaves similarly to dynamical friction and when the models are all normalised to one today, they present an increase in the growth factor. These considerations explain the behaviour found in Fig. 1. We also note that changing $\Gamma$ has no effect on the evolution of perturbations, but this is only true for the $\Lambda$CDM model, where density, and hence pressure, do not evolve in time. Finally, the effects induced by dissipative pressure friction are much larger than for dynamical friction. This happens because whilst the contribution of dynamical friction is constant in time, dissipative friction changes with time and its effects are stronger later in the cosmic evolution.
Similar considerations can be made for Figs. 2 and 3 where we show the effects of dynamical friction for the first and dissipative friction for the latter for the six dark energy models investigated in this work. We note that the effect of dynamical friction for these models is qualitatively the same as for the $\Lambda$CDM model when we consider the relative difference with respect to a model without friction. This is clearly understood as dark energy is only affecting the background and the dynamical equation to be solved is the same for both $\Lambda$CDM and the dynamical dark energy models. Furthermore, the correction due to $\eta_{0}$ is constant in time.
A deeper discussion about the effect of dissipative pressure friction is worth. In this case, because pressure changes in time, the parameter $\Gamma$ has an effect and the higher it is (in absolute value), the higher is the difference with respect to the model not affected by friction. As for the $\Lambda$CDM model, a positive $\eta_{2}$ implies a higher growth factor. In addition, exactly because pressure changes in time, the observed effect depends on the particular equation of state. Models 1, 3, 4 and 5 share similar effects for the cases $(+1,-2)$ and $(+1,-1)$ while for model 2 differences start to increase. Model 6 presents the strongest effect for the combination $(+1,-2)$ and this is due to the fact that the equation of state stays in the phantom regime all the time and, as a consequence, pressure is higher (in absolute value).
V Evolution of nonlinear perturbations
The study of nonlinear perturbations is based on the approach of the spherical collapse, whose equations have been derived in Section II. We consider once again two perturbed fluids, dark matter and baryons, where only the first experiences the effects of dynamical friction. To derive the initial conditions of the full nonlinear equation we consider that the averaged matter density $\delta_{\rm m}$ diverges Pace et al. (2010). However, for numerical stability, we do not use $\delta_{\rm m}$, but rather its inverse, so that the collapse takes place when $1/\delta_{\rm m}\to 0$, as discussed in Pace et al. (2017b).
Some important quantities which can be defined within the formalism of the spherical collapse are the turn-around scale factor, $a_{\rm ta}$ and the nonlinear overdensity at turn-around $\zeta$. The first represents the time at which the spherical perturbation reaches its maximum extension before collapsing and shrinking, while the latter is defined as $\zeta=1+\delta_{\rm NL}(a_{\rm ta})$. The nonlinear overdensity at turn-around is constant for an Einstein-de Sitter model but for any other model, including the $\Lambda$CDM one, changes in time. We verified that dynamical friction has either a negligible or very small effect both on the turn-around scale factor $a_{\rm ta}$ or on the nonlinear overdensity at turn-around, $\zeta$. When considering dissipative pressure friction, instead, the value of $a_{\rm ta}$ changes up to $2\%$, while $\zeta$ differs up to $15\%$ with respect to a model without friction.
@@ V Evolution of nonlinear perturbations
Figure 4: Evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ for the $\Lambda$CDM model. In the top (bottom) panel we show results for the dynamical (dissipative pressure) friction. Line styles and colours are as in Fig. 1.
@@
Two additional and very important quantities, also from an observational point of view, are the linearly extrapolated overdensity $\delta_{\rm c}$ and the virial overdensity $\Delta_{\rm vir}$. The linear overdensity $\delta_{\rm c}$ is, together with the linear growth factor, the main ingredient of the halo mass function which, as the name says, represents the linear evolution of matter density perturbations, whose initial conditions are found to have the nonlinear matter density evolution to diverge at collapse time. The linear overdensity thus represents the solution of a linear differential equation but with initial conditions derived from its nonlinear counterpart. Its time evolution as a function of the collapse redshift $z$ for the $\Lambda$CDM model is presented in Fig. 4. We show the results for the dark energy models in Section A in Figs. 10 (dynamical friction) and 11 (dissipative pressure friction), to which we refer for the discussion of our findings.
@@
Even though the overdensity $\delta_{\rm c}$ is a linear quantity, the differential equation is evolved starting from nonlinear initial conditions and this explains why differences are enhanced with respect to the growth factor but are reduced with respect to the nonlinear overdensity $\zeta$. Focusing again on the $\Lambda$CDM model, for the dynamical friction case, we see that, as expected, differences are enhanced when the free parameter $\eta_{0}$ increases, so that for $\eta_{0}=0.01$ the linear overdensity is about 0.2% lower than when $\eta_{0}=0$, regardless of the collapse redshift considered. The highest difference is at $z=0$ and at the higher redshift all the models recover the corresponding value for the EdS model. When friction becomes stronger, $\delta_{\rm c}$ is lower, even though the initial conditions are enhanced. This is because to collapse at the same time of a model without friction, perturbations require a lower threshold for the collapse to compensate the presence of friction.
@@
In the bottom panel of Fig. 4 we consider dissipative pressure friction applied to the $\Lambda$CDM model. As expected, the effects induced by friction are higher than for dynamical friction and, in analogy to the growth factor, the parameter $\Gamma$ is irrelevant (but it will not be for dynamical dark energy models). Differences, now, are of the order of $-5\%$ and $\delta_{\rm c}$ decreases in agreement to what found above. This time, though, for all the combinations of parameters, we recover the standard $\Lambda$CDM model at higher redshift. This is expected as for dissipative pressure friction the pressure terms grows with time and is, therefore, important at relatively late times, when the dark energy component dominates.
@@
The virial overdensity $\Delta_{\rm vir}$ represents the value the nonlinear overdensity reaches at virialization. To determine how virialization proceeds, we follow Carrilho et al. (2022) for interacting dark energy models. Even though this is not the physical case we are studying here, for smooth dark energy, the equations of motion are the same and dark matter experiences, in both cases, an additional dynamical friction force.
@@
We can make explicit the different terms entering into the virial theorem. To do so, we follow the notation of Cataneo et al. (2019); Carrilho et al. (2022) and denote with [^7] We corrected a typo in the relation between $\delta$ and $y$ in Eq. (28) of Carrilho et al. (2022) and Eq. (34) of Cataneo et al. (2019).
@@
with $N\equiv\ln{a}$ where we decomposed the total potential energy $W$ in its contributions, i.e., $W=W_{\rm N}+W_{\rm de}+W_{\rm fr}$, with $W_{\rm N}$ the Newtonian potential due to matter perturbations, $W_{\rm de}$ the background dark energy contribution and $W_{\rm fr}$ the friction contribution. In a more familiar form, expressing these quantities in term of the physical radius $R$, we have
@@
Figure 5: Evolution of the virial overdensity $\Delta_{\rm vir}$ for the $\Lambda$CDM model. In the top (bottom) panel we show results for the dynamical (dissipative pressure) friction. Line styles and colours are as in Fig. 1.
Therefore, according to Eq. (38), to find the virilization epoch, we solve the equation
@@
with $a_{\rm c}$ and $a_{\rm vir}$ the collapse and virialization scale factor, respectively. The nonlinear overdensity $\delta_{\rm NL}$ is evaluated at the virialization epoch.
@@
Knowing the virialization epoch, using the formalism outlined above, we can also determine the ratio of the virial radius to the turn-around radius, $R_{\rm vir}/R_{\rm ta}$. This quantity gives an idea of how much the perturbation shrinks from its maximum size once it reaches the virial equilibrium. This quantity is, therefore, affected by friction but it is also cosmology- and time-dependent. It is constant in time only for an EdS model, for which $R_{\rm vir}/R_{\rm ta}=1/2$ (in absence of friction contributions). Once again, for dynamical friction differences with respect to the standard case are extremely small, but can reach a few percent in the case of dissipative pressure friction.
@@
We can now discuss the effects of friction on the evolution of the virial overdensity $\Delta_{\rm vir}$, as defined in Eq. (44) and presented in Fig. 5 for the $\Lambda$CDM model. In the top (bottom) panel we consider dynamical (dissipative pressure) friction. Once more, the first case leads to very small effects, except for $\eta_{0}=0.01$, with differences of about $0.5\%$, relatively constant in time, with a small increase at late times when the cosmological constant dominates. For the case of dissipative pressure friction, effects are, as extensively discussed before, relevant, reaching up to $50\%$ today. These results are in perfect agreement with our previous analysis of the different nonlinear quantities, in particular of the overdensity at the turn-around $\zeta$.
@@
Concluding, dynamical friction appears having only small signatures on the models we investigated. On the other side, albeit theoretically-motivated, the dynamical pressure model shows highly-performing changes into the shapes that are expected once friction is involved. Thus, on the one hand, the latter clearly appears more probable to be measured. Conversely, the first case appears hard to be disentangled from the background dark energy. Remarkably, both are not excluded, since it would be possible to experimentally constrain the associated free parameters to check which values are thus expected to correctly reproduce a friction contribution. For the sake of completeness, we expect the dynamical pressure model to behave less deeply, being constructed in the simplest way in which the pressure itself is modified as due to the inclusion of a friction term.
VI Effects of the modified evolution of the mass function

### v2

@@ 4 Evolution of linear perturbations
Although small, $\eta$ affects the initial conditions of Eq. (5), similarly to what happens in early dark energy and massive neutrinos scenarios. This happens because the friction term $\eta/H$ is small but not negligible. To elucidate this concept, let us denote by $\eta_{\rm ini}=\eta(a_{\rm ini})/H(a_{\rm ini})\ll 1$ the value to be used at an initial time to evaluate the initial conditions. In general, at an early time, a given cosmological model is well approximated by an EdS model, so that we can assume $\Omega_{\rm m}(t)\approx 1$ and $H^{\prime}/H\approx-3/2$ and that the growth factor follows a power-law behavior, $\delta=Aa^{n}$, where $A$ is an arbitrary constant amplitude that can be safely set to one. Plugging the Ansatz for $\delta$ into Eq. (5), we need to solve a second order algebraic equation for the exponent $n$:
@@
whose growing mode solution is, for $\eta_{\rm ini}\ll 1$, $n_{+}=1-2\eta_{\rm ini}/5$ if curvature and radiation are neglected. This result is expected from physical considerations: friction slows down the collapse and $n_{+}<1$ when $\eta\neq 0$. We can compare this result with a generic early dark energy model. If we denote by $\Omega_{\rm e}\ll 1$ the amount of early dark energy, we have $\Omega_{\rm m}=1-\Omega_{\rm e}$ and find for the growing mode $n_{+}=1-3\Omega_{\rm e}/5$. This same expression applies to massive neutrinos, if we replace $\Omega_{\rm e}$ with the fraction of massive neutrinos $f_{\nu}=\Omega_{\nu}/\Omega_{\rm m}$, with $\Omega_{\nu}$ the neutrinos density parameter. Note the similarity of the different physical systems. In this case, too, therefore, structure formation is slowed down. This means that to match the growth rate of structures in the $\Lambda$CDM model, the initial amplitude of the perturbations must be higher. We will discuss this in more detail when investigating the evolution of nonlinear perturbations.
After this discussion, we can now study the evolution of the growth factor. We consider two clustering species, baryons (denoted with the subscript $\mathrm{b}$) and dark matter (denoted with the subscript $\mathrm{cdm}$), of which only the latter is affected by dynamical friction. The two equations to be solved are
@@
$\delta_{\rm m}=\frac{\Omega_{\rm b}\delta_{\rm b}+\Omega_{\rm cdm}\delta_{\rm cdm}}{\Omega_{\rm b}+\Omega_{\rm cdm}}\,.$ (26)
4.1 Initial settings
If we consider a single clustering species, we can assume that at early times $\delta=A\,a^{n}$, where both $A$ and $n$ are constant. This leads to a second-order algebraic equation for $n$ whose solution, for $\eta=\eta_{0}H$ and $\eta_{0}\ll 1$, is $n=1-\tfrac{2}{5}\eta_{0}$. We can proceed in a similar way if we consider both dark matter and baryons. In this case, we can also assume a power law such that $\delta_{\rm b}=A\,a^{n}$ and $\delta_{\rm cdm}=B\,a^{n}$, with constant $A$, $B$, and $n$ to determine. We assume that the two perturbations have the same slope but different amplitudes. If we plug this Ansatz in the previous equations, we obtain, taking into account that at early times the background is extremely well represented by an EdS model,
@@ 4.1 Initial settings
$\frac{\delta_{\rm b}}{\delta_{\rm cdm}}=1+\frac{\eta_{\rm ini}}{n+1/2}\,,$ (28)
which shows that the baryon perturbation amplitude is increased with respect to that of cold dark matter, as baryons are not affected by friction. Also note that for vanishing friction, $\delta_{\rm b}=\delta_{\rm cdm}=\delta_{\rm m}$, as expected.
Plugging Eq. (28) into the equation for $\delta_{\rm cdm}$, leads to an equation for the exponent $n$:
@@
Note that when $a\ll 1$, $\Omega_{\rm b}=\Omega_{\rm b,0}/\Omega_{\rm m,0}$, with $\Omega_{\rm m,0}=\Omega_{\rm b,0}+\Omega_{\rm cdm,0}$. Hence, the term which involves baryons is always constant and subdominant with respect to unity. For realistic values of both $\Omega_{\rm b,0}$ and $\Omega_{\rm m,0}$, we have $\Omega_{\rm b}$ of the order of 0.2.
Although it is possible to rewrite it as a third-order equation, it is simpler to solve it as is, to avoid going through complex radicals. However, we can use the cubic equation to investigate the behavior of the initial slope when $\eta_{\rm ini}\ll 1$. The equation to be solved is
@@
If $\eta_{\rm ini}=0$, it is easy to see that the solutions are $n=-\tfrac{1}{2}$, $n=-\tfrac{3}{2}$ and $n=1$, of which only the latter two are the physical ones, corresponding to the decaying and growing modes, respectively. At first order in $\eta_{\rm ini}$, we obtain $n=1-\tfrac{2}{5}\eta_{\rm ini}(1-\Omega_{\rm b})$, where it is understood that $\Omega_{\rm b}$ is evaluated at early times. The term in parentheses is positive for realistic values of the baryon density, and hence, also in this case, we can see that friction dampens structure formation. If we set, for simplicity $1-\Omega_{\rm b}\approx 1$, we recover the previous result, showing the excellent agreement between the two approaches.
@@
Since the two equations of motion are different, we cannot combine them into a single equation for the variable $\delta_{\rm m}$. Therefore, to highlight the importance of friction and the differences induced by having two species rather than one, we consider the evolution of the linear growth factor $D_{+}$ as a function of the redshift $z$ considering either one or two species. As differences among the two different cases are of the order of $0.1\%$, for simplicity, we will consider just the case where both dark matter and baryons are affected by dynamical friction. This will also simplify the discussion when considering clustering dark energy models.
To justify the Ansatz previously used for the time evolution of baryons and cold dark matter, let us consider the two equations describing the evolution of perturbations in the standard case, when no friction contribution is present. In this case, the two equations are the same, and considering the definition of $\delta_{\rm m}$, the combined equation has the same form as the previous two, but with the term $\Omega_{\rm b}\delta_{\rm b}+\Omega_{\rm cdm}\delta_{\rm cdm}$ replaced by $\Omega_{\rm m}\delta_{\rm m}$. Hence, it is clear that the two variables evolve in the same way but have different amplitudes. As the friction term is very small, we can assume that this Ansatz is still valid in our case. This is certainly true for the equations that govern the growth factor, but similar considerations can be made later when discussing the equation for the spherical collapse model.
The qualitative discussion above can be made more quantitative following [90]. If we consider the difference $\delta_{\rm cb}=\delta_{\rm cdm}-\delta_{\rm b}$, this quantity decays as $a^{-1/2}$ in the EdS regime, showing that the two quantities will evolve in the same way and we can thus assume that they only differ in their amplitude when setting the initial conditions. For a more detailed treatment, we refer the reader to [91, 92].
4.2 Growth factor
@@
In this Section we present the effects of friction on the evolution of the linear growth factor, using the two recipes described above. We first consider its effect on the $\Lambda$CDM reference model. We then compare it with six different dark energy models, two of them characterized by a constant equation of state, $w_{\rm de}=-0.9$ and $w_{\rm de}=-1.1$ (DE1 and DE2), and four with a linearly time-dependent equation of state, $w_{\rm de}=w_{0}+w_{a}(1-a)$, with $w_{0}$ and $w_{a}$ constant [19, 20]. We choose the following combinations of parameters $(w_{0},w_{a})$: $(-0.9,\pm 0.3)$ (DE3 and DE4) and $(-1.1,\pm 0.3)$ (DE5 and DE6) to cover quintessence, phantom and barrier crossing models. [^8] These are the same parameters adopted for the DEMNUni simulations [93], in which we tested the implementation of our spherical collapse code in case of absence of friction contributions to show a very good quantitative agreement between the theoretical and numerical halo mass function.
Figure 1: Evolution of the growth factor for the $\Lambda$CDM model for the two recipes used in this work to describe cosmic friction. Left panel: dynamical friction as parametrised in Eq. (13). Right panel: dissipative pressure friction as parametrised in Eq. (14). Different colours and line styles refer to different values of the free parameters.
@@
In Fig. 1 we present the evolution of the linear growth factor, divided by the scale factor, for the $\Lambda$CDM model, showing results for both dynamical friction (left panel) and dissipative pressure friction (right panel). The considerations which we are going to make for this model will also apply to the other models studied in this work.
@@
As $\eta$ opposes the formation of structures, perturbations grow more slowly with respect to a model where the friction term is absent. Hence, because we want the same rate today (all the curves have been normalised to unity today), structures need to grow faster in the past. This explains the differences observed. Note that this behavior is analogous to $\Lambda$CDM versus EdS. Therefore, the higher the value of $\eta$, the higher the difference from a model without friction.
@@
For dynamical friction, in general, differences are very small and only appreciable for the highest value chosen $\eta_{0}=10^{-2}$. The effects of dissipative pressure friction follow the same trend. Dissipative pressure friction behaves similarly to dynamical friction, and when the models are all normalized to one today, they present an increase in the growth factor. These considerations explain the behavior found in Fig. 1. We also note that changing $\Gamma$ has no effect on the evolution of perturbations, but this is only true for the $\Lambda$CDM model, where density and hence pressure do not evolve in time. Finally, the effects induced by dissipative pressure friction are much larger than those induced by dynamical friction. This happens because whilst the contribution of dynamical friction is constant in time, dissipative friction changes with time and its effects are stronger later in the cosmic evolution.
@@
Figure 2: Evolution of the growth factor for the model DE1 for dynamical friction (left panel) and dissipative pressure friction (right panel). Line styles and colours are as in Fig. 1.
Similar considerations can be made for Fig. 2 where we show the effects of dynamical and dissipative friction for the model DE1. We do not show the other models because the qualitative behavior is the same and they only differ quantitatively. This applies to all the quantities we will present for dark energy models. The effect of dynamical friction for these models is similar to the $\Lambda$CDM model when we consider the relative difference with respect to a model without friction. This is clearly understood, as dark energy only affects the background and the dynamical equation to be solved is the same for both $\Lambda$CDM and the dynamical dark energy models. Furthermore, the correction due to $\eta_{0}$ is constant in time.
A deeper discussion on the effect of dissipative pressure friction is worth pursuing. In this case, because pressure changes over time, the parameter $\Gamma$ has an effect and the higher it is (in absolute value), the higher the difference with respect to the model not affected by friction. For the $\Lambda$CDM model, a positive $\eta_{2}$ implies a higher growth factor. In addition, because pressure changes over time, the observed effect depends on the particular equation of state. Models DE1, DE3, DE4 and DE5 share similar effects for the cases $(+1,-2)$ and $(+1,-1)$ while for model DE2 differences start to increase. Model DE6 presents the strongest effect for the combination $(+1,-2)$ and this is due to the fact that the equation of state remains in the phantom regime all the time and, as a consequence, the pressure is higher (in absolute value).
5 Evolution of nonlinear perturbations
The study of nonlinear perturbations is based on the approach of the spherical collapse, whose equations have been derived in Sect. 2. We consider once again two perturbed fluids, dark matter and baryons, where only the first experiences the effects of dynamical friction. To derive the initial conditions of the full nonlinear equation we consider that the averaged matter density $\delta_{\rm m}$ diverges [41]. However, for numerical stability, we do not use $\delta_{\rm m}$, but rather its inverse, so the collapse takes place when $1/\delta_{\rm m}\to 0$, as discussed in [62].
Some important quantities which can be defined within the formalism of spherical collapse are the turn-around scale factor, $a_{\rm ta}$ and the nonlinear overdensity at turn-around $\zeta$. The first represents the time at which the spherical perturbation reaches its maximum extension before collapsing and shrinking, while the latter is defined as $\zeta=1+\delta_{\rm NL}(a_{\rm ta})$. The nonlinear overdensity at turn-around is constant for an Einstein-de Sitter model, but for any other model, including the $\Lambda$CDM one, changes in time. We verified that dynamical friction has either a negligible or very small effect on both the turn-around scale factor $a_{\rm ta}$ and the nonlinear overdensity at turn-around, $\zeta$. When considering dissipative pressure friction, instead, the value of $a_{\rm ta}$ changes up to $2\%$, while $\zeta$ differs up to $15\%$ with respect to a model without friction.
Figure 3: Evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ for the $\Lambda$CDM model. In the left (right) panel we show results for the dynamical (dissipative pressure) friction. Line styles and colours are as in Fig. 1.
@@ 5 Evolution of nonlinear perturbations
Two additional and very important quantities, also from an observational point of view, are the linearly extrapolated overdensity $\delta_{\rm c}$ and the virial overdensity $\Delta_{\rm vir}$. Linear overdensity $\delta_{\rm c}$ is, together with the linear growth factor, the main ingredient of the halo mass function which, as the name says, represents the linear evolution of matter density perturbations, whose initial conditions are found to have the nonlinear matter density evolution to diverge at collapse time. The linear overdensity thus represents the solution of a linear differential equation but with initial conditions derived from its nonlinear counterpart. Its time evolution as a function of the collapse redshift $z$ for the $\Lambda$CDM model is presented in Fig. 3. We show the results for the dark energy models in A in Fig. 8, to which we refer for a discussion of our findings.
@@
Although the overdensity $\delta_{\rm c}$ is a linear quantity, the differences between the models are larger than with respect to the growth factor, despite the fact that we solve the same differential equation. This happens because for the growth factor the initial conditions are the same for all the models, while for $\delta_{\rm c}$ they are model dependent and are obtained by first solving a nonlinear differential equation. Focusing again on the $\Lambda$CDM model, for the dynamical friction case, we see that, as expected, differences are enhanced when the free parameter $\eta_{0}$ increases, so that for $\eta_{0}=0.01$ the linear overdensity is about 0.2% lower than when $\eta_{0}=0$, regardless of the collapse redshift considered. The highest difference is at $z=0$ and at a higher redshift all models recover the corresponding value for the EdS model. When friction becomes stronger, $\delta_{\rm c}$ is lower, even though the initial overdensity is larger. This is because to collapse at the same time as a model without friction, perturbations require a lower threshold for the collapse to compensate the presence of friction.
@@
In the right panel of Fig. 3 we consider dissipative pressure friction applied to the $\Lambda$CDM model. As expected, the effects induced by friction are higher than for dynamical friction and, analogously to the growth factor, the parameter $\Gamma$ is irrelevant (but it will not be for dynamical dark energy models). The differences, now, are of the order of $-5\%$ and $\delta_{\rm c}$ decreases, in agreement with what was found above. However, this time, for all combinations of parameters, we recover the standard $\Lambda$CDM model at higher redshift. This is expected as for dissipative pressure friction the pressure term grows with time and is, therefore, important at relatively late times, when the dark energy component dominates.
@@
The virial overdensity $\Delta_{\rm vir}$ represents the value that the nonlinear overdensity reaches at virialization. To determine how virialization proceeds, we follow the same procedure as in [57] for interacting dark energy models. Even though this is not the physical case we are studying here, for smooth dark energy the equations of motion are the same, and dark matter experiences, in both cases, an additional dynamical friction force.
@@
We can make explicit the different terms entering into the virial theorem. To do so, we follow the notation of [94, 57] and denote with [^9] We corrected a typo in the relation between $\delta$ and $y$ in Eq. (28) of [57] and Eq. (34) of [94].
@@
with $N\equiv\ln{a}$, and where we decomposed the total potential energy $W$ in its contributions, i.e., $W=W_{\rm N}+W_{\rm de}+W_{\rm fr}$, with $W_{\rm N}$ the Newtonian potential due to matter perturbations, $W_{\rm de}$ the background dark energy contribution and $W_{\rm fr}$ the friction contribution. In a more familiar form, expressing these quantities in term of the physical radius $R$, we have
@@
Figure 4: Evolution of the virial overdensity $\Delta_{\rm vir}$ for the $\Lambda$CDM model. In the left (right) panel we show results for the dynamical (dissipative pressure) friction. Line styles and colours are as in Fig. 1.
Therefore, according to Eq. (31), to find the virialization epoch, we solve the equation
@@
with $a_{\rm c}$ and $a_{\rm vir}$ the collapse and virialization scale factor, respectively. The nonlinear overdensity $\delta_{\rm NL}$ is evaluated in the virialization epoch.
@@
Knowing the virialization epoch, using the formalism outlined above, we can also determine the ratio of the virial radius to the turn-around radius, $R_{\rm vir}/R_{\rm ta}$. This quantity gives an idea of how much the perturbation shrinks from its maximum size once it reaches the virial equilibrium. This quantity is, therefore, affected by friction, but it is also cosmology- and time-dependent. It is constant in time only for an EdS model, for which $R_{\rm vir}/R_{\rm ta}=1/2$ (in the absence of friction contributions). Once again, the effect due to dynamical friction is extremely small, but it is more relevant, about a few percent for dissipative pressure friction.
@@
We can now discuss the effects of friction on the evolution of virial overdensity $\Delta_{\rm vir}$, as defined in Eq. (37) and presented in Fig. 4 for the $\Lambda$CDM model. In the left (right) panel we consider dynamical (dissipative pressure) friction. Once more, the first case leads to very small effects, except for $\eta_{0}=10^{-2}$, with differences of about $0.5\%$, relatively constant in time, with a small increase at late times when the cosmological constant dominates. For the case of dissipative pressure friction, the effects are relevant, as extensively discussed before, reaching up to $50\%$ today. These results are in perfect agreement with our previous analysis of the different nonlinear quantities, in particular the overdensity at the turn-around $\zeta$.
@@
In conclusion, dynamical friction appears to have only small signatures in the models we investigated. On the other side, albeit theoretically motivated, the dynamical pressure model shows highly-performing changes into the shapes that are expected once friction is involved. Thus, on the one hand, the latter clearly appears more probable to be measured. Conversely, the first case appears to be hard to disentangle from the background dark energy. Remarkably, both are not excluded since it would be possible to experimentally constrain the associated free parameters to check which values are thus expected to correctly reproduce a friction contribution. For the sake of completeness, we expect the dynamical pressure model to behave less deeply, being constructed in the simplest way in which the pressure itself is modified as a result of the inclusion of a friction term.
6 Effects of the modified evolution of the mass function